# Self-Adaptive Graph Mixture of Models

### v1

@@ Appendix A Case Study
The effectiveness of the SAGMM method is demonstrated through a meticulously designed scenario below. As shown in (xu2018powerful), GNNs can be at most as powerful as the WL test where the input to the GNN is considered as $\boldsymbol{1}$ or $C\boldsymbol{1}$, where $C$ is the graph operator. For the SAGE (hamilton2017inductive) model the graph operator, $C_{\text{SAGE}}=I_{n}+D^{-1}A$, with $D$ representing the degree matrices of the graph.
In this setup, we consider three experts: GCN (kipf2016semi), SAGE, and GIN (xu2018powerful) — under all one’s input as that of the WL test. Among these experts, GIN emerges as the preferred choice, where the discriminatory power of GIN is shown to be as powerful as the WL-test (xu2018representation). We demonstrate a gating function for the SAGMM model that effectively selects the GIN model under all one’s input for the first layer. The outputs of these experts at the first layer are given by: $H_{\text{GIN}}^{(1)}=\text{MLP}^{(1)}\left((A+(1+\epsilon^{(l)})I_{n})\boldsymbol{1}\right),\;H_{\text{GCN}}^{(1)}=\sigma\left(F^{-1}\boldsymbol{1}\Theta^{(0)}\right),\;H_{\text{SAGE}}^{(1)}=\sigma\left(2\boldsymbol{1}\Theta^{(0)}\right).$ $F$ is a diagonal matrix with entries $F_{ii}=\sum_{j}C_{\text{GCN}}(ij)$ The gating function for SAGMM is detailed in Algorithm 1 in main paper, but for brevity, let us consider a gating score function defined for an expert $e_{i}\in\{\text{GCN, SAGE, GIN}\}$ and batch $b$ as
@@
In Eq. (18), Var denotes the variance operation. It can be observed that the gating scores for GCN, SAGE will be zero. The ideal architecture in this case, GIN will be selected for which the variance will be non-zero. The advantage of the SAGMM model described in the current scenario is distinct from prior works on a mixture of experts, such as (wang2023graphmixtureexpertslearning; yao2024moe), where the model relies on a single model type and applies MoE within intermediate layers. The selected model type in these prior works must be GIN to achieve optimal performance.
@@ B.1 Background
Matrix multiplication plays a fundamental role in machine learning and scientific computing, with numerous optimization techniques explored in prior works such as (pytorchamm), (guennebaud2010eigen), and (abadi2016tensorflow). Count sketch, a powerful dimensionality reduction method introduced in (Count-sketch) maps an $n$-dimensional vector $u$ to a lower-dimensional space of size $c$. This transformation utilizes a random hash function $h:[n]\rightarrow[c]$ and a Rademacher variable $s:[n]\rightarrow\{-1,1\}$. The dimensionality reduction process is defined as
@@
Tensor sketch extends the concept of count sketch (Count-sketch) and is widely used for efficient dimensionality reduction in large-scale machine learning applications (Tensorsketchfirst). Pham and Pag (Phamandpag) introduced an efficient implementation leveraging the Fast Fourier Transform (FFT) and its inverse, formulated as
@@
In (ding2022sketchgnn), tensor sketching is applied to approximate the element-wise $k$-th power of a matrix product:
@@
To study the impact of pruning experts during training, we investigate the loss of the GNN under diverse mixtures of experts. Towards this, we present Theorem 1 (as listed in main paper also), where we investigate the GNN with a mixture of experts considering randomized inputs to the GNN. The input feature matrix to the GNN considered are $X=[x_{1},x_{2},\ldots,x_{n}]$, where each feature $x_{u}$ is independently sampled from a multivariate normal distribution, $\mathcal{N}(\mathbf{0}_{d},I_{d\times d})$. Let the experts be given by $\{e_{1},e_{2}\ldots e_{N}\}$. The investigation of GNN properties using such randomized inputs is inspired by (kanatsoulis2022representation). The article demonstrates how randomized inputs for the input feature matrix are a more effective approach for analyzing GNN behaviour compared to the inputs used in the WL test.
@@ B.2 Assumptions of Theorem 1
The proof builds upon the approximation of GNN update rules in Eq. (1) as introduced in (ding2022sketchgnn). The non-linearity in the update rule is approximated using a power series expansion. Tensor-Sketch (tensorsketch) is then applied to approximate the series. For a single convolution matrix $C^{(l)}$ with expansion order $r$ we have the following,
@@
where the matrix $A_{1e_{i}u}\in\mathbb{R}^{1\times n}=c_{1e_{i}}\text{TS}_{1}(C_{e_{i}u,:}^{(0)})({R^{(1)}})^{T}$. Here, $R^{(1)}\in\mathbb{R}^{n\times c}$ is the count-sketch matrix  (tensorsketch), with $c$ representing a dimension smaller than $n$.
@@
Applying the Reverse Markov inequality (abhishek2019introduction) to the random variable $\mathcal{L}(\frac{\eta(\mathcal{Y}_{u})}{k_{u}})$, we obtain:
@@
By applying Jensen’s inequality (boyd2004convex), since $\mathcal{L}$ is a convex function of $\mathcal{Y}_{u}$ (as log-sum-exp, -log is convex), and using the fact that $\mathbb{E}(\mathcal{Y}_{u})=0$, we obtain:
@@ C.1 Datasets Description
  • Deezer-Europe (rozemberczki2020characteristic): It is a user-user network of European users of the Deezer music streaming platform. Nodes represent individual users, and edges denote mutual friendships between them. Node features are derived from the artists liked by these users. The task is to classify users’ gender. Following the benchmark setup in (lim2021newbenchmarkslearningnonhomophilous), the dataset is split randomly into 50%/25%/25% for training, validation, and testing.
  • YelpChi (mukherjee2013yelp): This dataset is a collection of user reviews from Yelp for businesses (hotels and restaurants) located in the Chicago area. The dataset contains both filtered (suspected fake) and unfiltered (genuine) reviews as identified by Yelp’s internal spam filter. The task is to detect deceptive reviews using a binary classification setting, where filtered reviews are treated as fake and unfiltered ones as non-fake . The dataset follows the real-world class imbalance naturally present in Yelp’s data.
  • ogbn-proteins: This is a graph where nodes represent proteins, and edges indicate biologically significant associations, categorized by type based on species. The dataset contains proteins from 8 species. The task is a multi-label binary classification to predict the presence or absence of 112 protein functions. The public OGB (hu2020open) split is used.
  • ogbn-arxiv: This dataset is a citation network of Computer Science (CS) papers on arxiv. Nodes represent papers, and edges indicate citation relationships. Each node is associated with a 128-dimensional feature vector derived by averaging the word embeddings of the paper’s title and abstract, generated using the WORD2VEC model. The objective is to predict the subject area of each paper, covering 40 distinct subject areas. The dataset follows the split defined in (hu2020open).
  • Pokec (jure2014snap): It is a large-scale social network dataset containing various profile features, such as geographical region, registration time, and age. The task is to predict users’ gender based on these features. The dataset is split randomly into 50% for training, 25% for validation, and 25% for testing.
  • ogbn-products: Itis based on the Amazon Co-Purchasing network (mcauley2015inferringnetworkssubstitutablecomplementary), where nodes represent products, and edges signify frequent co-purchases between products. Node features are created using bag-of-words representations of product descriptions, and the labels correspond to the top-level categories of the products. The dataset follows public OGB split.
  • ogbn-papers100M: This dataset is a directed citation graph of 111 million papers indexed by MAG (10.1162/qss_a_00021). The task is to predict the subject areas of the subset of papers that are published in arXiv. The dataset follows public OGB split.
@@
  • ogbg-mol* (Molecular Graphs): The ogbg-molhiv and ogbg-molpcba datasets are two molecular property prediction datasets of different sizes: ogbg-molhiv (small) and ogbg-molpcba (medium). Each graph represents a molecule, where nodes are atoms, and edges are chemical bonds. Input node features are 9-dimensional, containing atomic number and chirality, as well as other additional atom features such as formal charge and whether the atom is in the ring. Input edge features are 3-dimensional, containing bond type, bond stereochemistry as well as an additional bond feature indicating whether the bond is conjugated. The task is to predict the target molecular properties as accurately as possible, where the molecular properties are cast as binary labels. The dataset follows OGB (hu2020open) split.
@@ C.3 Experiment Settings
For larger datasets, where the graph sizes prevent full-graph training, we utilize a mini-batch training approach similar to the method in (wu2023nodeformerscalablegraphstructure). At the beginning of each training epoch, the nodes are shuffled and partitioned into random mini-batches of size B. During each iteration, one mini-batch is selected, and the corresponding subgraph is extracted from the original graph. This subgraph is fed into the model to compute the loss on the training nodes within that batch. For ogbn-products and ogbn-papers100M we employ neighbor sampling training approach.
@@ Appendix D SAGMM Architecture Components
  • Gating Mechanism: The gating network dynamically routes each node to a subset of experts using a thresholded attention mechanism based on SGA (wu2023sgformer). Node representations are first transformed using learnable weight matrices and then compared via scaled dot-product similarity. The resulting attention logits are processed using a sigmoid gating function and a learnable threshold vector. A ReLU activation is then applied to retain only positive scores, thereby selecting the most relevant experts per node based on the gated activations.

### v2

@@ Appendix A Case Study
The effectiveness of the SAGMM method is demonstrated through a meticulously designed scenario below. As shown in (Xu et al. 2018a), GNNs can be at most as powerful as the WL test where the input to the GNN is considered as $\boldsymbol{1}$ or $C\boldsymbol{1}$, where $C$ is the graph operator. For the SAGE (Hamilton et al. 2017a) model the graph operator, $C_{\text{SAGE}}=I_{n}+D^{-1}A$, with $D$ representing the degree matrices of the graph.
In this setup, we consider three experts: GCN (Kipf and Welling 2016), SAGE, and GIN (Xu et al. 2018a) — under all one’s input as that of the WL test. Among these experts, GIN emerges as the preferred choice, where the discriminatory power of GIN is shown to be as powerful as the WL-test (Xu et al. 2018b). We demonstrate a gating function for the SAGMM model that effectively selects the GIN model under all one’s input for the first layer. The outputs of these experts at the first layer are given by: $H_{\text{GIN}}^{(1)}=\text{MLP}^{(1)}\left((A+(1+\epsilon^{(l)})I_{n})\boldsymbol{1}\right),\;H_{\text{GCN}}^{(1)}=\sigma\left(F^{-1}\boldsymbol{1}\Theta^{(0)}\right),\;H_{\text{SAGE}}^{(1)}=\sigma\left(2\boldsymbol{1}\Theta^{(0)}\right).$ $F$ is a diagonal matrix with entries $F_{ii}=\sum_{j}C_{\text{GCN}}(ij)$ The gating function for SAGMM is detailed in Algorithm 1 in main paper, but for brevity, let us consider a gating score function defined for an expert $e_{i}\in\{\text{GCN, SAGE, GIN}\}$ and batch $b$ as
@@
In Eq. (18), Var denotes the variance operation. It can be observed that the gating scores for GCN, SAGE will be zero. The ideal architecture in this case, GIN will be selected for which the variance will be non-zero. The advantage of the SAGMM model described in the current scenario is distinct from prior works on a mixture of experts, such as (Wang et al. 2023; Yao et al. 2024), where the model relies on a single model type and applies MoE within intermediate layers. The selected model type in these prior works must be GIN to achieve optimal performance.
@@ B.1 Background
Matrix multiplication plays a fundamental role in machine learning and scientific computing, with numerous optimization techniques explored in prior works such as (Paszke et al. 2017), (Guennebaud et al. 2010), and (Abadi et al. 2016). Count sketch, a powerful dimensionality reduction method introduced in (Charikar et al. 2002) maps an $n$-dimensional vector $u$ to a lower-dimensional space of size $c$. This transformation utilizes a random hash function $h:[n]\rightarrow[c]$ and a Rademacher variable $s:[n]\rightarrow\{-1,1\}$. The dimensionality reduction process is defined as
@@
Tensor sketch extends the concept of count sketch (Charikar et al. 2002) and is widely used for efficient dimensionality reduction in large-scale machine learning applications (Pham and Pagh 2013c). Pham and Pag (Pham and Pagh 2013a) introduced an efficient implementation leveraging the Fast Fourier Transform (FFT) and its inverse, formulated as
@@
In (Ding et al. 2022), tensor sketching is applied to approximate the element-wise $k$-th power of a matrix product:
@@
To study the impact of pruning experts during training, we investigate the loss of the GNN under diverse mixtures of experts. Towards this, we present Theorem 1 (as listed in main paper also), where we investigate the GNN with a mixture of experts considering randomized inputs to the GNN. The input feature matrix to the GNN considered are $X=[x_{1},x_{2},\ldots,x_{n}]$, where each feature $x_{u}$ is independently sampled from a multivariate normal distribution, $\mathcal{N}(\mathbf{0}_{d},I_{d\times d})$. Let the experts be given by $\{e_{1},e_{2}\ldots e_{N}\}$. The investigation of GNN properties using such randomized inputs is inspired by (Kanatsoulis and Ribeiro 2022). The article demonstrates how randomized inputs for the input feature matrix are a more effective approach for analyzing GNN behaviour compared to the inputs used in the WL test.
@@ B.2 Assumptions of Theorem 1
The proof builds upon the approximation of GNN update rules in Eq. (1) as introduced in (Ding et al. 2022). The non-linearity in the update rule is approximated using a power series expansion. Tensor-Sketch (Pham and Pagh 2013b) is then applied to approximate the series. For a single convolution matrix $C^{(l)}$ with expansion order $r$ we have the following,
@@
where the matrix $A_{1e_{i}u}\in\mathbb{R}^{1\times n}=c_{1e_{i}}\text{TS}_{1}(C_{e_{i}u,:}^{(0)})({R^{(1)}})^{T}$. Here, $R^{(1)}\in\mathbb{R}^{n\times c}$ is the count-sketch matrix  (Pham and Pagh 2013b), with $c$ representing a dimension smaller than $n$.
@@
Applying the Reverse Markov inequality (Abhishek et al. 2019) to the random variable $\mathcal{L}(\frac{\eta(\mathcal{Y}_{u})}{k_{u}})$, we obtain:
@@
By applying Jensen’s inequality (Boyd and Vandenberghe 2004), since $\mathcal{L}$ is a convex function of $\mathcal{Y}_{u}$ (as log-sum-exp, -log is convex), and using the fact that $\mathbb{E}(\mathcal{Y}_{u})=0$, we obtain:
@@ C.1 Datasets Description
  • Deezer-Europe (Rozemberczki and Sarkar 2020): It is a user-user network of European users of the Deezer music streaming platform. Nodes represent individual users, and edges denote mutual friendships between them. Node features are derived from the artists liked by these users. The task is to classify users’ gender. Following the benchmark setup in (Lim et al. 2021), the dataset is split randomly into 50%/25%/25% for training, validation, and testing.
  • YelpChi (Mukherjee et al. 2013): This dataset is a collection of user reviews from Yelp for businesses (hotels and restaurants) located in the Chicago area. The dataset contains both filtered (suspected fake) and unfiltered (genuine) reviews as identified by Yelp’s internal spam filter. The task is to detect deceptive reviews using a binary classification setting, where filtered reviews are treated as fake and unfiltered ones as non-fake . The dataset follows the real-world class imbalance naturally present in Yelp’s data.
  • ogbn-proteins: This is a graph where nodes represent proteins, and edges indicate biologically significant associations, categorized by type based on species. The dataset contains proteins from 8 species. The task is a multi-label binary classification to predict the presence or absence of 112 protein functions. The public OGB (Hu et al. 2020) split is used.
  • ogbn-arxiv: This dataset is a citation network of Computer Science (CS) papers on arxiv. Nodes represent papers, and edges indicate citation relationships. Each node is associated with a 128-dimensional feature vector derived by averaging the word embeddings of the paper’s title and abstract, generated using the WORD2VEC model. The objective is to predict the subject area of each paper, covering 40 distinct subject areas. The dataset follows the split defined in (Hu et al. 2020).
  • Pokec (Jure 2014): It is a large-scale social network dataset containing various profile features, such as geographical region, registration time, and age. The task is to predict users’ gender based on these features. The dataset is split randomly into 50% for training, 25% for validation, and 25% for testing.
  • ogbn-products: Itis based on the Amazon Co-Purchasing network (McAuley et al. 2015), where nodes represent products, and edges signify frequent co-purchases between products. Node features are created using bag-of-words representations of product descriptions, and the labels correspond to the top-level categories of the products. The dataset follows public OGB split.
  • ogbn-papers100M: This dataset is a directed citation graph of 111 million papers indexed by MAG (Wang et al. 2020b). The task is to predict the subject areas of the subset of papers that are published in arXiv. The dataset follows public OGB split.
@@
  • ogbg-mol* (Molecular Graphs): The ogbg-molhiv and ogbg-molpcba datasets are two molecular property prediction datasets of different sizes: ogbg-molhiv (small) and ogbg-molpcba (medium). Each graph represents a molecule, where nodes are atoms, and edges are chemical bonds. Input node features are 9-dimensional, containing atomic number and chirality, as well as other additional atom features such as formal charge and whether the atom is in the ring. Input edge features are 3-dimensional, containing bond type, bond stereochemistry as well as an additional bond feature indicating whether the bond is conjugated. The task is to predict the target molecular properties as accurately as possible, where the molecular properties are cast as binary labels. The dataset follows OGB (Hu et al. 2020) split.
@@ C.3 Experiment Settings
For larger datasets, where the graph sizes prevent full-graph training, we utilize a mini-batch training approach similar to the method in (Wu et al. 2023a). At the beginning of each training epoch, the nodes are shuffled and partitioned into random mini-batches of size B. During each iteration, one mini-batch is selected, and the corresponding subgraph is extracted from the original graph. This subgraph is fed into the model to compute the loss on the training nodes within that batch. For ogbn-products and ogbn-papers100M we employ neighbor sampling training approach.
@@ Appendix D SAGMM Architecture Components
  • Gating Mechanism: The gating network dynamically routes each node to a subset of experts using a thresholded attention mechanism based on SGA (Wu et al. 2023b). Node representations are first transformed using learnable weight matrices and then compared via scaled dot-product similarity. The resulting attention logits are processed using a sigmoid gating function and a learnable threshold vector. A ReLU activation is then applied to retain only positive scores, thereby selecting the most relevant experts per node based on the gated activations.